OpenStudy (anonymous):

Solve: x^2 - 7x = -12

OpenStudy (anonymous):

use middle term factor process

OpenStudy (anonymous):

Got it! =5 and 4

OpenStudy (anonymous):

no, 3 and 4 check it out again

OpenStudy (pfenn1):

\[x^2-7x+12=0\]Can either factor or might be just as easy to use the quadratic equation.\[ax^2+bx+c=0\]\[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

apply the concept correctly

Parth (parthkohli):

Add 12 to both sides. \(\Large \color{MidnightBlue}{\Rightarrow x^2 - 7x + 12 }\) A quadratic equation is ax^2 + bx + c = 0 According to that, the formula is: \(\Large \color{MidnightBlue}{\Rightarrow x = {-b \pm \sqrt{b^2 - 4ab} \over 2a} }\)

OpenStudy (anonymous):

Here's a nother problem i'm not sure on... Quadratic equations can be solved using the quadratic formula... Always, sometimes, or never.. I think its sometimes

OpenStudy (anonymous):

x= 3 and 4... http://www.wolframalpha.com/input/?i=x%5E2+-+7x+%3D+-12

Parth (parthkohli):

Always...they can be....unless there's no solution

OpenStudy (anonymous):

always there is solution.. no unless is there thing is that the roots may be real or imaginary depending on the equation

Parth (parthkohli):

No solution here means no real solution. There may be a complex one.

OpenStudy (anonymous):

but if a particular set is specified, then the answer will be 'sometimes'

OpenStudy (anonymous):

What about this one... What are the solutions of x^2 + 2x + 9

OpenStudy (anonymous):

i got the answer x = -2 + or - √32 all over 2

Parth (parthkohli):

There's always the quadratic formula :D

OpenStudy (anonymous):

use the quadratic formula again...... simple

Parth (parthkohli):

You can look at wolframalpha to check it.

OpenStudy (anonymous):

i got the answer x = -2 + or - √32 all over 2